OpenStudy (anonymous):

Can you check my work please? derivative of 2tan(x/2)-5

OpenStudy (anonymous):

by letting u = (x/2) y=f(u)=2tan(u) dy/du=(2)(sec^2u) du/dx=(1) y=2sec^(u) y' = 2sec^2(x/2) ?

OpenStudy (mimi_x3):

not quite

OpenStudy (anonymous):

yeah... because the answer is only sec^2(x/2)... can you tell me where i went wrong?

OpenStudy (mimi_x3):

you forgot to cancel out the 2

OpenStudy (anonymous):

the 2? mmm, you mean the 2 and the x/2?

OpenStudy (mimi_x3):

wait a sec

OpenStudy (anonymous):

ok

OpenStudy (mimi_x3):

\[ u = \frac{x}{2} = \frac{du}{dx} = \frac{1}{2}\] \[y = 2 tan(u) - 5 => \frac{dy}{dx} = 2sec^{2}(x/2) * \frac{1}{2}\] \[ \frac{dy}{dx} = sec^{2}(x/2)\]

OpenStudy (mimi_x3):

ok this is some chain rule which i cant remmeber

OpenStudy (mimi_x3):

@Miracrown: what do you think?

Miracrown (miracrown):

mim, i think a lot :P

OpenStudy (anonymous):

ahhh yes yes. now i get it. :)

OpenStudy (anonymous):

thanks :D gonna post another problem.

OpenStudy (mimi_x3):

:)